# The PLATO Mission

### v1

@@ 3.1.2 Expected PLATO Planet Detection Yield
Up to now, the expected yield of planet detections has been studied in the ESA Definition Study Report (hereafter called "Red Book", ESA-SCI 109), by [147, 207], and by forthcoming publications like Cabrera et al. (in prep.). We summarise these results here as the status at the current stage of the mission development and for comparison with future PLATO yield estimates by the community. For completion, there are also studies on the fraction of planets that will be impacted by background contaminants which are also valuable for properly determining the scope of follow-up efforts [269, 50].
@@
The total planet detection yield depends on the number of target fields observed as well as the duration of continuous observations per field. The baseline for ESA operations of PLATO is an observing period with a total duration of 4 years. However, mission extensions are anticipated, subject to ESA review. The satellite is designed with consumables for 8.5 years of operations in total, permitting significant mission extensions. The goal to detect planets in the HZ of solar-like stars drives the observing strategy of PLATO and requires a minimum of two years continuous observation per field (with longer periods preferred). However, shorter periods have also been considered to estimate the full planet yield potential to study different kinds of systems. In the Red Book and in Heller et al [147], a baseline observing scenario of two long instrument pointings of two years each ("2+2" scenario) is compared to a "3+1" scenario. The latter allows for one year with shorter observations of, e.g., 60 days duration each. Cabrera et al. consider also extended observation periods to study the benefits for long-period planet detection (see also below).
@@
For the Red Book, planet occurrence rates were taken from Fressin et al [114] based on Kepler data for all planets, except for small planets in the HZ which were not well constrained. To cover the wide range of uncertainty for HZ-planets (see Table 1), we assumed a planet occurrence rate of 40% as baseline, but also computed for extreme values such as 2% and 100%. Heller et al [147] assumed planet occurrence rates spanning from 37% to 88% [from 53] for earths in the HZ. Cabrera et al. use the same assumptions as in the Red Book for consistency, but also study additional scenarios not shown here. Matuszewski et al [207] use a very different approach. They address this unknown factor by using planet population models [107, 108] to predict the number of planets formed up to the HZ.
Transit search algorithms
@@
The efficiency of transit signal detection algorithms depends on several factors [see, e.g., 79]. It is relatively straightforward to quantify how the transit detection efficiency depends on the strength of the signal searched for. The signal depends on the astrophysical system (planet to star radius ratio, orbital parameters, etc.) and on the level of noise in the light curve (photon noise and systematic – red – noise). The noise budget is determined by the instrument design (e.g. aperture diameter, number of cameras observing the same target) and the target magnitude (see Section 11). In addition, variability in the light curves plays a crucial role, whether it is of instrumental origin (e.g. effects like random telegraphic pixels, tearing effect, non-stability of temperature or voltages, etc.) or resulting from the stellar properties (e.g. intrinsic variability, binarity). This variability can be partly corrected by light curve processing methods (e.g. for most systematic trends). The variability fraction that cannot be removed remains as residual noise budget in the processed light curves.
@@
At the time of writing the Red Book, detailed instrument performance studies for PLATO as shown in Section 11 were not available yet. The best existing analogy were Kepler data in combination with the available transit detection pipeline. It was therefore decided to account for all the above effects by choosing global detection efficiency factors based on the experience from Kepler [114] and assuming that the PLATO mission noise requirements will be reached. Accordingly, for the P1 sample 100% detection efficiency was assumed for planets down to Earth size, based on the fact that targets are sufficiently bright and photon noise dominates over other noise sources for P1. This fraction (100%) was justified by i) estimating a signal-to-noise ratio of 13 (respectively 15) observing 2 (respectively 3) consecutive transits with depth 84 ppm (actually, with Sun-like limb darkening values, the transit depth of a planet with the size of the Earth is closer to 120 ppm, see e.g. Heller 146) and ii) considering the detectability fraction originally in Jenkins et al [164] that is discussed in Fressin et al [114]. Note that 34 ppm is the worst value of any P1 star, not the average noise. The median of the noise distribution (which was uncertain at that time) can be estimated around 20 ppm, producing a signal-to-noise ratio of 21. Note also that Christiansen et al [78] would have revised the 100% fraction to 90%. However, this has a limited impact in the yield for Earth-sized planets in 1 year orbit when comparing to the uncertainty in the occurrence rate (ranging from 2% to 100%, see Table 1). Residual stellar intrinsic noise could still impact the detection efficiency for the smallest planets, but we anticipated that the well-chosen target stars in P1 would not be dominated by this effect (by choosing quiet stars). For large planets (Neptune-sized planets and larger) in the fainter P5 sample, it is appropriate to assume 100% signal detection efficiency as well. Planets with 4 Earth radii produce a signal-to-noise ratio of 20 with just 1 transit around stars observed with 240 ppm, representing the worst of the P5 distribution (magnitude 13 observed at the edge of the field of view). Our assumed detection pipeline efficiency for P5 is reduced to 50% for planets with radii REarth in light curves with 80 ppm in 1 hour or better. Here the signal-to-noise ratio is 7 for 3 consecutive transits, which corresponds to 50% fraction as per Jenkins et al [164], but see the discussion above. Again, here we took the worst noise as dimensioning case, which is conservative. For a P5 star with 50 ppm the signal-to-noise ratio with 2 transits is 9. Light curves with higher residual noise levels are not considered for our estimate of small planet yields, although larger planets would still be detectable.
@@
Heller et al [147] have improved this approach by applying their own transit signal detection algorithm [Transit Least Squares 148] to transits inserted into simulated PLATO light curves created by the PLATO Solar-like Light-curve Simulator [PSLS 268]. This approach mimicked the response to the known instrumental noise sources of PLATO in a somewhat more realistic manner. Matuszewski et al [207] and Cabrera et al. (in prep.) use an updated noise budget based on the currently known instrumental noise sources of PLATO and information from the recent PIC which was not yet available when the Heller et al [147] study was performed.
@@
Table 2 shows the resulting planetary transit event yields. For comparison, we summarise in column 2 the current knowledge of confirmed transiting planets in the literature. These planets result from ground-based detections as well as CoRoT, Kepler, TESS, and CHEOPS observations. Today, fewer than 1200 confirmed transiting planets around stars brighter than mag are known, and none of them is a small planet in the HZ of a solar-like star. Fewer than 400 planets orbit stars brighter than mag and are therefore suitable for effective RV follow-up.
@@
For host stars of to 13 mag the resulting planet properties from PLATO will be similar to the main part of the planet populations provided by the Kepler mission, hence masses will be available only for objects of special interest justifying dedicated follow-up campaigns (i.e. as for CoRoT-7b or Kepler-10b of mag) or from Transit Time Variations (TTV) analysis. Nevertheless, these planets around fainter stars will be highly useful to study e.g. radius-orbit distributions, etc., just as for similar population analyses carried out with Kepler data. Accounting for an improved age calibration of stars from PLATO also for classical methods, we may be able to go beyond Kepler and study, e.g., the radius gap and orbital distance distributions as a function of stellar age also for this faint population. Considering planets of all sizes and all orbital periods, about 4600 planets (see Table 2) can be detected by PLATO around stars brighter than mag by combining the P1 sample with the P5 sample. The PLATO planet yield in this magnitude range is therefore similar to the total planet yield in Kepler, but most Kepler planets orbit stars of 14 mag and are consequently difficult to characterise. Thus, at the bright end, where they are more suitable for further characterisation, the number of expected PLATO detections is higher than what has been achieved with Kepler. Indeed, although the somewhat smaller effective aperture of PLATO leads to higher noise levels at the same magnitude (see Section 11), this is compensated by the significantly wider field-of-view of PLATO. This is a consequence of PLATO’s modular camera approach and design.
@@
The estimates by Heller et al [147], as well as those by Cabrera et al. using more realistic stellar distributions and noise budget levels, are within the uncertainty ranges presented in the Red Book. The high end of numbers from Matuszewski et al [207] stem from using a planet formation model for the assumed planet occurrence rate. The example shows the need for observational constraints to planet formation models.
@@
The key planet sample of PLATO consists of small planets in the HZ of solar-like stars. The Red Book estimates a wide uncertainty range (6–280) for the yield of such planets around bright (11 mag) stars, resulting from the large uncertainty in planet occurrence rates. These planets will mostly be suitable for RV follow-up, eventually depending on telescope time available, and form the main data product from PLATO for exoplanetary science. The study by Heller et al [147] predicts a smaller number of planets in the HZ, but that is because these authors concentrated exclusively on Earth-sized planets ( R) orbiting P1 stars, while the Red Book estimates consider planets of R and stars with mag in the P5 sample too. Cabrera et al. (in prep.) agree with previous results with a large uncertainty range, which is again associated with different assumptions for planet occurrence rates. Matuszewski et al [207] use a planet synthesis model to estimate the expected planet occurrences. This approach results in much higher numbers of predicted planets, all other assumptions being the same. Figure 3 illustrates the expectations for PLATO for small planets in the HZ of solar-like stars for the 2+2 scenario addressed in the Red Book with current knowledge of the end-of-life (EOL) instrument performance of PLATO (see Section 11) using PIC version 1.1.0. The figure shows roughly 50 stars in the HZ which is in agreement with the prediction range of 0 to 95 in Table 2, performing simulations as described above.
@@
In summary, at this point the predictions on the number of planets to be detected by PLATO are highly dependent on the respective assumed planet occurrence rates, in particular for temperate small planets. Hence, these predictions merely reflect our poor knowledge on planet formation efficiencies for small planets. Improving this situation with PLATO determined planet occurrence rates from a homogeneously analysed sample is a key science driver for PLATO.
@@
To illustrate the expected impact from PLATO further, we compare PLATO with TESS choosing the validated TESS Objects of Interest (TOI) of the 2 year nominal mission [135]. Figure 4 shows a comparison of TOIs orbiting FGK stars compared to PLATO expectations for a single field observed for 2 years. The predicted number of planets orbiting K stars detected by PLATO is only modestly increased compared to that expected from TESS. For G and F type stars, however, PLATO is expected to outnumber TESS results by factors 2-3 (G stars: 571/TESS and 1595/PLATO; F stars: 468/TESS and 1448/PLATO). It is interesting to compare the performance on temperate small planets of both missions for the same 2-year observing duration. Most TESS TOIs have orbits 10 days and only very few have orbits 27 days. PLATO planets will show a significantly larger fraction of longer periods from 27 days to 100 days. This is also the case for planets orbiting K stars, even though the total number of planets is comparable. Figure 4 shows how the PLATO efficiency to detect small planets around K, G, and F stars largely outnumbers TESS’s performance when considering the same observing period.
@@
The expected planet yield for the TESS extended mission has been addressed in Kunimoto et al [183], showing a factor 3 increased number of expected new detections for a total of 7 years of operations. This includes those to be discovered from analyses of the full-frame images, as well as the higher cadence of 2-minute TESS light curves. However, also for the TESS extended mission most planets are expected to have orbits up to about 30 days and will be larger than 4 R (their Fig. 6). Regarding small temperate planets in the HZ of solar-like stars (see Fig. 3) we compare PLATO’s capacity with Fig. 5 in Kunimoto et al [183]: in their bin of 2 R and orbital periods between 250 and 400 days, no planet detections are foreseen. This illustrates that we have to await the results from PLATO to fill this parameter range, which includes the range Earth would inhabit. It is the larger aperture (see also Section 11) and the observing strategy of PLATO that will allow breakthroughs in this parameter space.
@@
Figure 5 shows a prediction of the total number of planets discovered by PLATO as a function of the pointing duration. This prediction is done for a single pointing field (Cabrera et al. in prep.). Note that only planets with orbital periods up to 418 days were simulated [114]. Beyond this value we have no statistics, so longer-period planets will not add to the values in our analysis. Nevertheless, even if their transit probability is small, theoretical models predict that such long-period planets could be numerous [e.g. 207]. Figure 5 shows that hot Jupiter detections saturate very quickly, within one year of observation duration. In this saturated regime, observing 2 independent fields for 2 years each would double the amount of planets detected, compared with the yield from one field targeted for 4 years. We note, however, that by the time of PLATO launch, the TESS mission already significantly increased our knowledge on these kind of hot gas giants and they may not be the main driver for selecting PLATO fields and their observing mode. The number of detected hot super earths saturates within 2 to 4 years of observations. Hence, once saturation is reached, additional detections can only be gained when changing the target field. In contrast to the hot planets, the number of temperate earths grows almost linearly with the pointing duration for up to 10 years. We clearly see the benefit in the detection yield of temperate planets when increasing the observing duration of a target field. This behaviour points towards increasing the observation of PLATO´s first target field beyond the nominal 2 years. However, before finally deciding on extended observing durations, we also have to consider the availability of resources for follow-up and the real in-flight instrument performance that might affect the expected detection efficiency. Once these factors are better known, a refined assessment on PLATO´s detection efficiencies as a function of target field observing duration can be made.
As outlined in Rauer et al [254], we expect additional planet detections around, e.g., binary (multiple) stars (including circum-binary planets), subgiant and giant stars, as well as the potential for exo-comets, planetary moons, rings or Trojan planets. Estimates for the efficiency of PLATO detections on these targets are subject for future studies and beyond the scope of this review. This applies also for planet detections made via, e.g., Transiting Time Variations (TTV) analyses or reflected light observations. Here, we only recall these extra methods and targets as a reference for the many new discoveries expected from the PLATO mission.

### v2

@@ 3.1.2 Expected PLATO Planet Detection Yield
Up to now, the expected yield of planet detections has been studied in the ESA Definition Study Report (hereafter called "Red Book", ESA-SCI 118), by [160, 229], and by forthcoming publications like Cabrera et al. (in prep.). We summarise these results here as the status at the current stage of the mission development and for comparison with future PLATO yield estimates by the community. For completion, there are also studies on the fraction of planets that will be impacted by background contaminants which are also valuable for properly determining the scope of follow-up efforts [292, 54].
@@
The total planet detection yield depends on the number of target fields observed as well as the duration of continuous observations per field. The baseline for ESA operations of PLATO is an observing period with a total duration of 4 years. However, mission extensions are anticipated, subject to ESA review. The satellite is designed with consumables for 8.5 years of operations in total, permitting significant mission extensions. The goal to detect planets in the HZ of solar-like stars drives the observing strategy of PLATO and requires a minimum of two years (taking the Earth-Sun system as baseline) continuous observation per field (with longer periods preferred). However, shorter periods have also been considered to estimate the full planet yield potential to study different kinds of systems. In the Red Book and in Heller et al [160], a baseline observing scenario of two long instrument pointings of two years each ("2+2" scenario) is compared to a "3+1" scenario. The latter allows for one year with shorter observations of, e.g., 60 days duration each. Cabrera et al., in prep., consider also extended observation periods to study the benefits for long-period planet detection (see also below).
@@
For the Red Book, planet occurrence rates were taken from Fressin et al [125] based on Kepler data for all planets, except for small planets in the HZ which were not well constrained. To cover the wide range of uncertainty for HZ-planets (see Table 1), we assumed a planet occurrence rate of 40% as baseline, but also computed for extreme values such as 2% and 100%. Heller et al [160] assumed planet occurrence rates spanning from 37% to 88% [from 58] for earths in the HZ. Cabrera et al. use the same assumptions as in the Red Book for consistency, but also study additional scenarios not shown here. Matuszewski et al [229] use a very different approach. They address this unknown factor by using planet population models [116, 117] to predict the number of planets formed up to the HZ.
Transit detection efficiency
@@
The efficiency of transit signal detection algorithms depends on several factors [see, e.g., 83]. It is relatively straightforward to quantify how the transit detection efficiency depends on the signal strength. The signal depends on the astrophysical system (planet to star radius ratio, orbital parameters, etc.) and on the level of noise in the light curve (photon noise and systematic – red – noise). This noise budget is determined by the instrument design (e.g. aperture diameter, number of cameras observing the same target) and the target magnitude (see Section 11). In addition, temporal variability in the light curves plays a crucial role, whether it is of instrumental origin (e.g. effects like random telegraphic pixels, tearing effect, non-stability of temperature or voltages, etc.) or resulting from the stellar properties (e.g. intrinsic variability, binarity). This variability can be partly removed by light curve processing methods (e.g. for most systematic trends). Remaining variability is left in the processed light curves as residual noise.
@@
At the time of writing the Red Book in 2017, detailed instrument performance studies for PLATO as shown in Section 11 were not available. The best existing analogy were Kepler data. It was therefore decided to account for all the above noise effects by choosing global detection efficiency factors based on the experience from Kepler [125] and assuming the PLATO mission noise requirements will be reached. We studied the P1 and P5 samples with the following assumptions: P1 sample: 100% detection efficiency for planets down to Earth size (photon noise dominates over other noise sources for P1). This was justified by i) estimating a signal-to-noise ratio of 13 (respectively 15) observing 2 (respectively 3) consecutive transits with depth 84 ppm (actually, with Sun-like limb darkening values, the transit depth of a planet with the size of the Earth is closer to 120 ppm, see e.g. Heller 159) and ii) considering the detectability fraction originally in Jenkins et al [175] that is discussed in Fressin et al [125].
Residual stellar intrinsic noise could still impact the detection efficiency for the smallest planets, but we anticipated the well-chosen target stars in P1 not to be dominated by this effect (hence choosing quiet stars). P5 sample: For large planets (Neptune-sized and larger) 100% signal detection efficiency was assumed. Planets with 4 Earth radii produce a signal-to-noise ratio of 20 with just 1 transit around stars observed with 240 ppm, representing the worst of the P5 distribution (magnitude 13 observed at the edge of the field of view). The detection efficiency for P5 targets reduces to 50% for planets with radii 2 REarth in light curves with 80 ppm in 1 hour or better. Here the signal-to-noise ratio is 7 for 3 consecutive transits, which corresponds to 50% fraction as per Jenkins et al [175]. Again, we took a conservative case. For a P5 star with 50 ppm the signal-to-noise ratio with 2 transits is 9. Light curves with higher residual noise levels are not considered for our estimate of small planet yields, although larger planets would still be detectable.
@@
Heller et al [160] have improved this approach by applying their own transit signal detection algorithm [Transit Least Squares 162] to transits inserted into simulated PLATO light curves created by the PLATO Solar-like Light-curve Simulator [PSLS 291]. This approach mimicked the response to the known instrumental noise sources of PLATO in a somewhat more realistic manner. Matuszewski et al [229] and Cabrera et al. (in prep.) use an updated noise budget based on the currently known instrumental noise sources of PLATO and information from the recent PIC which was not yet available when the Heller et al [160] study was performed.
@@
Table 2 shows the resulting planetary transit event yields using the various approaches and input catalogues/input population as discussed above. For comparison, we summarise in column 2 the current knowledge of confirmed transiting planets in the literature. These planets result from ground-based detections as well as CoRoT, Kepler, TESS, and CHEOPS observations. Today, fewer than 1500 confirmed transiting planets around stars brighter than mag are known, and none of them is a small planet in the HZ of a solar-like star. Fewer than 500 planets orbit stars brighter than mag and are therefore suitable for effective RV follow-up.
@@
For host stars of to 13 mag the resulting planet properties from PLATO will be similar to the main part of the planet populations provided by the Kepler mission, hence masses for small, terrestrial planets will be available only for objects of special interest justifying dedicated follow-up campaigns (i.e. as for CoRoT-7b or Kepler-10b of mag) or from Transit Time Variations (TTV) analysis. Nevertheless, these planets around fainter stars will be highly useful to study e.g. radius-orbit distributions, etc., just as for similar population analyses carried out with Kepler data. Accounting for an improved age calibration of stars from PLATO also for classical methods, we may be able to go beyond Kepler and study, e.g., the radius gap and orbital distance distributions as a function of stellar age also for this faint population. Considering planets of all sizes and all orbital periods, about 4600 planets (see Table 2) can be detected by PLATO around stars brighter than mag by combining the P1 sample with the P5 sample. The PLATO planet yield in this magnitude range is therefore similar to the total planet yield in Kepler, but most Kepler planets orbit stars of 14 mag and are consequently difficult to characterise. Thus, at the bright end, where they are more suitable for further characterisation, the number of expected PLATO detections is higher than what has been achieved with Kepler. Indeed, although the somewhat smaller effective aperture of PLATO leads to higher noise levels at the same magnitude (see Section 11), this is compensated by the significantly wider field-of-view of PLATO. This is a consequence of PLATO’s modular camera approach and design.
@@
The estimates by Heller et al [160], as well as those by Cabrera et al. using more realistic stellar distributions and noise budget levels, are within the uncertainty ranges presented in the Red Book. The high end of numbers from Matuszewski et al [229] stem from using a planet formation model for the assumed planet occurrence rate. The formation model produces higher planet occurences than assumed by the other authors based on, e.g., Kepler results. The example shows the need for observational constraints to planet formation models.
@@
The key planet sample of PLATO consists of small planets in the HZ of solar-like stars. The Red Book estimates a wide uncertainty range (6–280) for the yield of such planets around bright (11 mag) stars, resulting from the large uncertainty in planet occurrence rates. These planets will mostly be suitable for RV follow-up, eventually depending on telescope time available, and form the main data product from PLATO for exoplanetary science. The study by Heller et al [160] predicts a smaller number of planets in the HZ, but that is because these authors concentrated exclusively on Earth-sized planets ( 1.5 REarth) orbiting P1 stars, while the Red Book estimates consider planets of 2 REarth and stars with mag in the P5 sample too. Cabrera et al. (in prep.) agree with previous results with a large uncertainty range, which is again associated with different assumptions for planet occurrence rates. Matuszewski et al [229] use a planet synthesis model to estimate the expected planet occurrences. This approach results in much higher numbers of predicted planets, all other assumptions being the same. Figure 3 illustrates the expectations for PLATO for small planets in the HZ of solar-like stars for the 2+2 scenario addressed in the Red Book with current knowledge of the end-of-life (EOL) instrument performance of PLATO (see Section 11) using PIC version 1.1.0. The figure shows roughly 50 stars in the HZ which is in agreement with the prediction range of 0 to 95 in Table 2, performing simulations as described above.
@@
In summary, at this point the predictions on the number of planets to be detected by PLATO are highly dependent on the assumed planet occurrence rates, in particular for temperate small planets. Hence, these predictions merely reflect our poor knowledge on planet formation efficiencies for small planets. Improving this situation with PLATO determined planet occurrence rates from a homogeneously analysed sample is a key science driver for PLATO.
@@
To illustrate the expected impact from PLATO further, we compare PLATO with TESS choosing the validated TESS Objects of Interest (TOI) of the 2 year nominal mission [147]. Figure 4 shows a comparison of TOIs orbiting FGK stars compared to PLATO expectations for a single field observed for 2 years. The predicted number of planets orbiting K stars detected by PLATO is only modestly increased compared to that expected from TESS. For G and F type stars, however, PLATO is expected to outnumber TESS results by factors 2-3 (G stars: 571/TESS and 1595/PLATO; F stars: 468/TESS and 1448/PLATO). It is interesting to compare the performance for temperate small planets of both missions for the same 2-year observing duration. Most TESS TOIs have orbits 10 days and only very few have orbits 27 days. PLATO planets will show a significantly larger fraction of longer periods from 27 days to 100 days. This is also the case for planets orbiting K stars, even though the total number of planets is comparable. Figure 4 shows how the PLATO efficiency to detect small planets around K, G, and F stars largely outnumbers TESS’s performance when considering the same observing period.
@@
The expected planet yield for the TESS extended mission has been addressed in Kunimoto et al [198], showing a factor 3 increased number of expected new detections for a total of 7 years of operations. This includes those to be discovered from analyses of the full-frame images, as well as the higher cadence of 2-minute TESS light curves. However, also for the TESS extended mission most planets are expected to have orbits up to about 30 days and will be larger than 4 REarth (their Fig. 6). Regarding small temperate planets in the HZ of solar-like stars (see Fig. 3) we compare PLATO’s capacity with Fig. 5 in Kunimoto et al [198]: in their bin of 2 REarth and orbital periods between 250 and 400 days, no planet detections are foreseen. This illustrates that we have to await the results from PLATO to fill this parameter range, which includes the range Earth would inhabit. It is the larger aperture (see also Section 11) and the observing strategy of PLATO that will allow breakthroughs in this parameter space.
@@
Figure 5 shows a prediction of the total number of planets discovered by PLATO as a function of the pointing duration. This prediction is done for a single pointing field (Cabrera et al. in prep.). Note that only planets with orbital periods up to 418 days were simulated [125]. Beyond this value we have no statistics, so longer-period planets will not add to the values in our analysis. Nevertheless, even if their transit probability is small, theoretical models predict that such long-period planets could be numerous [e.g. 229]. Figure 5 shows that hot Jupiter detections saturate very quickly, within one year of observation duration. In this saturated regime, observing 2 independent fields for 2 years each would double the amount of planets detected, compared with the yield from one field targeted for 4 years. We note, however, that by the time of PLATO launch, the TESS mission already significantly increased our knowledge on these kind of hot gas giants and they may not be the main driver for selecting PLATO fields and their observing mode. The number of detected hot super earths saturates within 2 to 4 years of observations. Hence, once saturation is reached, additional detections can only be gained when changing the target field. In contrast to the hot planets, the number of temperate earths grows almost linearly with the pointing duration for up to 10 years. We clearly see the benefit in the detection yield of temperate planets when increasing the observing duration of a target field. This behaviour points towards increasing the observation of PLATO´s first target field beyond the nominal 2 years. However, before finally deciding on extended observing durations, we also have to consider the availability of resources for follow-up and the real in-flight instrument performance that might affect the expected detection efficiency. Once these factors are better known, a refined assessment on PLATO´s detection efficiencies as a function of target field observing duration can be made.
As outlined in Rauer et al [278], we expect additional planet detections around, e.g., binary (multiple) stars (including circumbinary planets), subgiant and giant stars, as well as the potential for exo-comets, planetary moons, rings or Trojan planets. Estimates for the efficiency of PLATO detections on these targets are subject for future studies and beyond the scope of this review. This applies also for planet detections made via, e.g., Transiting Time Variations (TTV) analyses or reflected light observations. Here, we only recall these extra methods and targets as a reference for the many new discoveries expected from the PLATO mission.